### v1

@@ 4 Empirical Findings
4.1 Impact of ChatGPT on User Activity in Stack Overflow
Our first research question investigates the impact of ChatGPT’s release on the content posted to Stack Overflow. Given the growing popularity of this LLM, we hypothesize that user engagement on the platform may decline over time, as ChatGPT could potentially address many of the queries posted there. We collected and categorized the relevant data to address the first research question (see Table 1). We then compared the frequency of questions, answers, comments, and tag usage before and after the introduction of ChatGPT. Notably, the number of submitted answers exhibits the most significant decline: 543 533543533543\,533543 533 answers before and 425 391425391425\,391425 391 answers after ChatGPT (i.e., -22%). This decline has a subsequent impact on the number of questions with accepted answers, which dropped from 212 775212775212\,775212 775 before ChatGPT to 157 187157187157\,187157 187 after ChatGPT (i.e., -26%). Comments and questions follow closely, showing a decrease of 19% and 14%, respectively. The same also applies to tag usage; when looking at the number of questions tagging Python, a decline of 22% was observed.
(a) Questions
(b) Answers
(c) Comments
Figure 4: Distribution of posted contents on Stack Overflow (questions, answers, and comments)
@@ 4.1 Impact of ChatGPT on User Activity in Stack Overflow
Upon close examination of Figures 3 and 4, depicting Stack Overflow content’s distribution and temporal evolution, trends emerge. Figure 3 reveals that before ChatGPT’s release, there was a higher density of posted content in the upper region of the violin plots. Post-release, this concentration appears more evenly distributed across the plots, including the middle and lower regions. Overall, the median daily count of submitted content is lower after ChatGPT’s introduction. Figure 4 shows the decline in posted content over time. Importantly, oscillating patterns present in the data before ChatGPT’s release continue to be observed (they correspond to weekends).
Next, we statistically assess whether there is a significant difference between the number of posted questions, answers, comments, and the usage of the five most adopted tags before and after ChatGPT’s release. Using the Wilcoxon rank sum test, we can reject the null hypothesis for all the evaluated contents (p-value <0.01). Such a result confirms our initial observational reported in Figures 3 and 4.
RQ1 answer: The introduction of ChatGPT has led to a statistically significant decline in user engagement on Stack Overflow. This is evidenced by the reduced frequency of posted questions, answers, comments, and tags usage, with the most pronounced decline observed in the number of submitted answers (22%). Furthermore, the data distribution visually supports the decline before and after ChatGPT’s release. Hence, ChatGPT’s introduction has impacted the platform’s dynamics, likely due to its ability to address many of the queries previously posted on Stack Overflow.
4.2 Comparing the Reliability of Answers: ChatGPT vs. LLaMA-2 vs Stack Overflow Users
For our second research question, we aim to assess the reliability of the answers generated by ChatGPT and LLaMA-2. Assuming that developers might use LLMs as assistive tools (Dakhel et al., 2023), it is important to ensure the information provided is wholly or partially reliable. To that end, we prompted the models to ask previously selected questions (i.e., 384 questions before and after ChatGPT’s release, i.e., 768 questions). Then, we checked the textual and semantic similarity between the generated and the accepted answer in Stack Overflow. Since users may consult or use ChatGPT and/or LLaMA-2 to answer questions in Stack Overflow, we focused on a specific time frame encompassing pre- and post-release periods of ChatGPT. This way, we might address concerns about potential shifts in answer quality after ChatGPT’s introduction, thus providing a comprehensive reliability evaluation. Figure 6 shows the distribution of the cosine metric across all prompted questions for both ChatGPT and LLaMA-2.
Textual Similarity Analysis
Regarding ChatGPT textual similarity, overall, both plots (i.e., before and after ChatGPT release) present similar densities, which means that there is no significant difference regarding the similarity between Stack Overflow and ChatGPT answers, before and after the ChatGPT’s introduction. The mean computed for each time frame validates this finding: 0.644 and 0.640, pre- and post-release, respectively. Our observation reported no significant statistical difference between both distributions (Wilcoxon, p-value = 0.59). This result indicates that ChatGPT is reliable enough to answer questions from Stack Overflow, while it is an expected result, considering that ChatGPT was trained on publicly available data until September 2021, eliminating possible associated bias (Carlini et al., 2022).
For LLaMA, the plots (i.e., before and after ChatGPT release) also present a similar density, observed by the average reported (0.616 and 0.619, before and after ChatGPT release), with a slightly more uniform similarity distribution for the “after ChatGPT release” plot. We also observed no significant statistical difference between both distributions (Wilcoxon test, p-value = 0.71), indicating that LLaMA-2 presented the same behavior for both sets of questions. Different from ChatGPT, this result might be unexpected considering that, though most of the data used to train LLaMA-2 were dated until September 2022, some tuning data was further required, dated up to July 2023, which might introduce bias in our results due to memorization (Carlini et al., 2022) [^11] https://github.com/facebookresearch/llama/blob/main/MODEL_CARD.md#training-data.
Since no significant difference was observed for the metric based on each LLM, aggregating the data for each LLM, we observe that LLaMA-2 presented an overall lower mean when compared to ChatGPT (0.617 and 0.642, respectively).A Wilcoxon test reports that ChatGPT was statically superior to LLaMA-2 (p-value <0.01) but with a small effect size (Cliff’s delta, 0.1). Hence, we conclude that though ChatGPT was statically superior, LLaMA-2 was close to achieving similar results, representing an alternative, free, and public option for the tech community.
Before ChatGPT’s release example. In general, as Figure 6 shows, ChatGPT and LLaMA-2 generate answers highly similar to those reported by Stack Overflow users. For example, consider the question with the highest similarity from the group of questions posted before ChatGPT’s release ($\theta$ = 0.92). In this question, a user asks How to get all keys whose values are null in Java 8 using Map [^12] https://stackoverflow.com/questions/73687017/. The user presents a HashMap with five keys representing colors associated with their HTML codes (two of which are null). Figure 5 presents the answers to this question in three versions: ① Figure 4(a) shows the answer generated by a human on Stack Overflow; ② Figure 4(b) shows the answer generated by ChatGPT; and ③ Figure 4(c) shows the answer generated by LLaMA-2.
@@
Both Stack Overflow and ChatGPT answers address the user’s request but adopt slightly different approaches. Overall, they adopt a stream object to process the HashMap elements; while the user works only with the set of keys (line 4), ChatGPT opts to work with the entire set (line 4). Next, to check whether the value associated with a key is null, the user first requests the value associated with a key to h and then checks whether the returned value is null by calling the method Objects.isNull (line 6). ChatGPT directly asks for the value associated with each set element, and for the applicable cases, it gets the associated keys (line 6). Finally, the selected keys are grouped similarly in a List for both cases (line 7).
@@
However, LLaMA-2’s solution does not work correctly ($\theta$ = 0.72). Indeed, the solution returns the values associated with the keys instead of just the keys themselves. It occurs because instead of initially getting the HashMap keys, the proposed solution gets the set of values (line 4). Then, for each value, it is checked whether they are null (line 6). For the valid cases, they are stored in the list used to save the results (line 7).
@@
After ChatGPT’s release example. Let us now consider the question with the highest similarity among the questions from the group of questions after ChatGPT’s release ($\theta$ = 0.92). In this example, the user asks for assistance regarding responsive design. [^13] https://stackoverflow.com/questions/75494817/ For that, the user provides some CSS code, defining different screen and font sizes, and based on screen size, it expects the elements to be adjusted. The Stack Overflow respondent provides four possible solutions for the problem. ChatGPT’s answer overlaps three: ① double-checking whether the expected HTML element is correctly targeting by the specified property; ② possible style conflicts overriding the expected font sizes; and ③ checking whether the browser window size is valid. Furthermore, the Stack Overflow respondent advises the user to use the !important declaration, ensuring that property is prioritized over others (style conflicts). ChatGPT adopts a more factual solution, asking the user to check whether the CSS file is correctly linked to the expected HTML file.
@@
LLaMA-2 takes a different route compared to ChatGPT and the original answer, by exploring solutions involving exclusively the code given as input ($\theta$ = 0.89). The solution proposed is related to possible properties being overridden by others, which ChatGPT and Stack Overflow answers also briefly discussed. To address the issue, LLaMA-2 suggests replacing the usage of the property max-width for min-width, or the previous two properties for just break-point. These examples show the power of ChatGPT and LLaMA-2 to correctly address different subjects, with a high level of reliability, compared to the answers provided by Stack Overflow users.
(a) ChatGPT
(b) LLaMA
@@
Evaluating reliability relying exclusively on textual similarity might introduce biases in our results. For example, when a user asks for instructions about exporting a fillable textbox PDF file using PowerBI, the accepted answer reports that the required action is not possible to be done [^14] https://stackoverflow.com/questions/74639660/export-pdf-with-fillable-textbox-from-powerbi. ChatGPT correctly informs that action is impossible and provides further alternatives that might support the user, resulting in a low textual similarity ($\theta$ = 0.10). This way, we may conclude that textual similarity represents a good metric for reporting high similarity between the answers, but not for reporting low similarity, as false negatives might take place. To address this threat, we also computed the semantic similarity between answers. Overall, LLaMA-2 reported a more positive evaluation with only 7.4% of the answers with low similarity, whereas ChatGPT reported a frequency of 22.1%. This result goes the opposite of our manual classification, where the frequency of low similarity was around 44% of our sample. Furthermore, we observe a low level of agreement between the LLMs’ evaluation, even for different runs of the same LLM.
@@
When computing the Krippendorff alpha between ChatGPT and LLaMA-2, we observe a coefficient of -0.02151515Adopting the same default temperature in both models.. This result shows that the LLMs’ ratings are not in good agreement, leading us to conclude that their ratings may not be reliable for the given data; furthermore, the negative value observed indicates less agreement than expected by chance. Next, to evaluate the agreement among different runs for the same model, we prompt ChatGPT to classify the answers three times. When computing the Krippendorff alpha for these runs, we observe better results but still a low coefficient (0.12). This result shows that the LLMs’, even on different runs of the same model, do not present consistent and stable results regarding semantic similarity. We also observe the same low similarity when evaluating the agreement of the LLMs with our manual analysis (Krippendorff alpha -0.02 and -0.01, LLaMA and ChatGPT, respectively). As a result, possible conclusions we draw here based on this semantic similarity would be biased by the random behavior of the LLMs (see Section 6).
@@
RQ2 answer: The reliability is statistically consistent. Both before and after ChatGPT’s release, the generated answers show a high degree of similarity to the accepted answers on Stack Overflow, as evidenced by the textual similarity (cosine metric). Although textual similarity seems to be a good metric for reporting answers with high similarity, this metric is not applicable for reporting low similarity, as false negatives might occur. Finally, the adopted semantic similarity metric based on LLMs classification also does not present an applicable metric for checking similarity among the answers.
@@
4.3 Challenges of LLMs in Generating Reliable and Accurate Answers Across Specific Domains
@@
This research question focuses on understanding the factors that might lead the LLMs to generate accurate answers and the associated challenges faced when generating these answers. Next, we discuss each aspect individually.
@@
To that end, first, we verify whether the context used to prompt the LLMs correlates with the textual similarity level previously computed (for that, we rely on the Spearman test). In our study, we consider the context as: ① the title of the question; ② the description of the question; and ③ the set of tags of the question. We have performed three distinct correlation analyses, one for each information. For the tags, we computed the number of tags associated with a question. In contrast, for the title and description, we first computed the number of words and then split them into intervals (2 and 25 words, respectively). For the title, we consider the interval of two words, as we did not observe questions with only a single word. For the description, we consider an interval of 25 words, as we consistently noted a proportional rise in the number of questions at each interval lenght (25, 50, etc). Eventually, for each question, we perform a Spearman test with each of these values and the question’ cosine similarity computed previously.
@@
From all performed analyses, the results show only a statistically significant correlation between the number of tags adopted for a question and the similarity for the answers generated post ChatGPT release (p-value <0.05). The correlation coefficient shows a negative relationship with $\rho=-0.11$, meaning that while one variable increases, the other might decrease. This result indicates that ChatGPT might be reliable in answering questions from specific and closed contexts. For example, the two questions discussed in Section 4.2 only adopt one single tag (java-8 and media-queries). However, this negative relationship reveals that ChatGPT might struggle when different contexts are present in the same question, as it has to arrange different topics to generate an answer. For instance, when a user asks for support regarding an unresolved reference, she uses three tags, which are all related (android, android-livedata, and android-lifecycle) [^16] https://stackoverflow.com/questions/75449324/unresolved-reference-to-transformations. For this question, ChatGPT generated an answer oppositely from the one provided by the Stack Overflow user, achieving a very low cosine similarity ($\theta$ = 0.06).
Sentiment Analysis
Regarding the sentiment analysis performed, we observe that most of the answers provided by Stack Overflow users and the LLMs present a neutral sentiment associated (96%, 97%, and 79%, Stack Overflow, ChatGPT, and LLaMA, respectively). Unlike other web forums, Stack Overflow users are used to keeping a formal, neutral, and direct way of speaking, which could be misinterpreted as a negative sign. When comparing the answers, we observe that ChatGPT reports a more neutral sentiment compared to the original answers from Stack Overflow users (Wilcoxon test, p-value <0.01). Furthermore, ChatGPT did not report any answer with a negative sentiment, while 13 user answers were classified with negative sentiment. Such a result reinforces the power of ChatGPT to adapt itself when answering questions, specifically in a tech context evaluated here.
Comparing LLaMA answers with those generated by ChatGPT and Stack Overflow users, we observe a significant statistical difference, revealing that LLaMA-2 report a less neutral sentiment (Wilcoxon test, p-value <0.01 and 0.01, respectively). However, unlike the previous answers, LLaMA-2 reports more answers with positive sentiment (20%). These results show that, in its current version, LLaMA-2 consistently adopts a positive-neutral sentiment when prompted by users, even in tech-related web forums.
Challenges for LLMs
Regarding the domain of the questions classified with low textual similarity for LLMs, we can make the same conclusions for both LLMs. Frameworks and Libraries represent the most challenging domain for the LLMs evaluated here (40% and 35%, ChatGPT and LLaMA, respectively), followed by Programming Languages (20% and 23%). When looking at the different sets of questions associated with each model, we observe they share 64% of the same questions, revealing that despite the structural differences between the models, they partially share the same limitations. Furthermore, less than 20% of all generated answers by both models presented a similarity of less than 0.5 (14% and 16%, for ChatGPT and LLaMA, respectively). This small frequency shows that even when the LLMs fail to create a more similar answer, they do not deviate from the original direction of the accepted answers.
On the other hand, for the questions with high textual similarity, the most promising domains are Programming Languages (40.5% and 40.3%), followed by Frameworks and Libraries (23.7% and 23%). Although the Frameworks and Libraries domain takes place in both analyses, we must discuss how the type of questions associated with each group plays an important role here. When users ask for specific, popular, and standard features of these dependencies, the LLMs achieve a high similarity; however, when the questions address general conventions or unpopular features about using these external dependencies, a low similarity is observed. For example, a user asks for support when failing to authenticate to an application using valid credentials. The solution accepted guides the user to hash the password when creating a new user in the application. [^17] https://stackoverflow.com/questions/75727923/authenticate-function-returning-none-in-login-api-view-django In its generated answer, ChatGPT even warns about the user’s credential and object creation steps without mentioning the actual cause.

### v2

@@ 4 Empirical Findings
4.1 How does the reliability of answers generated by ChatGPT and LLaMA-2 compare to those provided by Stack Overflow users?
In our first research question, after selecting Stack Overflow questions and prompting LLMs to address them, we aim to check their reliability based on textual (cosine metric) and semantic similarity.
(a) ChatGPT
(b) LLaMA
Textual Similarity Analysis
Regarding the textual similarity, Figure 3 shows the distribution of the cosine metric across all prompted questions for both ChatGPT and LLaMA-2. Specifically for ChatGPT, overall, both plots present similar densities (Figure 3(a), i.e., before and after ChatGPT release), indicating that there is no significant difference regarding the similarity between Stack Overflow and ChatGPT answers. By computing the mean for each time interval, we could validate our initial insight (0.644 and 0.640, pre- and post-release, respectively). Further comparing them, we report no significant statistical difference between both distributions (Wilcoxon, p-value = 0.59). By reporting such overall similar answers, ChatGPT shows its capability to address tech questions while adopting a similar behavior compared to humans at the textual level. Furthermore, we may highlight that such observation must consider that ChatGPT was trained on publicly available data until September 2021, eliminating possible associated bias. [^12] https://platform.openai.com/docs/models/gpt-3.5-turbo
For LLaMA, the plots also present a similar density, observed by the average reported (Figure 3(b), 0.616 and 0.619, before and after ChatGPT release, respectively), with a slightly more uniform similarity distribution for the “after ChatGPT release” plot. We also observed no significant statistical difference between both distributions (Wilcoxon test, p-value = 0.71), indicating that LLaMA-2 presented the same behavior for both sets of questions. Different from ChatGPT, this result might be unexpected considering that, though most of the data used to train LLaMA-2 were dated until September 2022, some tuning data was further required, dated up to July 2023, which might introduce bias in our results due to memorization [49] [^13] https://github.com/facebookresearch/llama/blob/main/MODEL_CARD.md#training-data.
Since no significant difference was observed for time interval metrics based on each LLM, we decided to compare the LLMs by aggregating their data. As a result, LLaMA-2 presented an overall lower mean when compared to ChatGPT (0.617 and 0.642, respectively). A Wilcoxon test reports that ChatGPT was statistically superior to LLaMA-2 (p-value <0.01) but with a small effect size (Cliff’s delta, 0.1). Hence, we conclude that though ChatGPT was statistically superior, LLaMA-2 was close to achieving similar results, representing an alternative, free, and public option for the tech community. To delve into the capabilities of the LLMs evaluated, we discuss some examples as follows.
Before ChatGPT’s release. Considering the question with the highest similarity by ChatGPT ($\theta$ = 0.92), the user places a question about a programming task in Java, asking about How to get all keys whose values are null in Java 8 using Map [^14] https://stackoverflow.com/questions/73687017/. An initial code snippet is provided to illustrate the task, presenting a HashMap with five keys representing colors associated with their HTML codes (two of which are null). Figures  4(a),  4(b), and  4(c) present the answers provided by a human on Stack Overflow, and generated by ChatGPT and LLaMA-2, respectively.
@@ 1 Map <String, String> h;
Both Stack Overflow and ChatGPT answers address the user’s request but adopt slightly different approaches. Initially, they adopt a stream object to process the HashMap elements; while the human answer works only with the set of keys (line 4), ChatGPT opts to work with the entire set (line 4). Next, to check whether the value associated with a key is null, the human answer first requests the value associated with a key to h and then checks whether the returned value is null by calling the method Objects.isNull (line 6). ChatGPT directly asks for the value associated with each set element, and for the applicable cases, it gets the associated keys (line 6). Finally, the selected keys are grouped similarly in a List for both cases (line 7).
@@
However, LLaMA-2’s solution does not work correctly ($\theta$ = 0.89). Indeed, the solution returns the values associated with the keys instead of just the keys themselves. It occurs because instead of initially getting the HashMap keys, the proposed solution gets the set of values (line 4). Then, for each value, it is checked whether they are null (line 6). For the valid cases, they are stored in the list used to save the results (line 7). Such a result shows that though LLaMA-2 was close to addressing the requested task, the reported answer requires further adjustments. An experienced user could run the code and detect the misbehavior. Next, with a few updates, the code could be fixed and behave as expected.
@@
After ChatGPT’s release. Now consider another question with the highest similarity achieved by ChatGPT dated after its release ($\theta$ = 0.92). In this example, the user asks for assistance regarding responsive design. [^15] https://stackoverflow.com/questions/75494817/ For that, the user provides some CSS code, defining different screen and font sizes, and based on screen size, it expects the elements to be adjusted. The Stack Overflow respondent provides four possible solutions for the problem: [^16] https://stackoverflow.com/a/75494849/5141439
{quoting}
① Make sure that the .my-element class is being applied to the correct element in your HTML. ② Check that there are no other styles elsewhere in your CSS that might be overriding the font size changes made by the media queries. ③ Try adding the !important declaration to the font-size property in each media query. ④ Verify that your browser window size is within the range specified.
ChatGPT’s answer overlaps three: ① double-checking whether the expected HTML element is correctly targeting by the specified property; ② possible style conflicts overriding the expected font sizes; and ③ checking whether the browser window size is valid. Furthermore, the Stack Overflow respondent advises the user to use the !important declaration, ensuring that property is prioritized over others (style conflicts). ChatGPT adopts a more factual solution, asking the user to check whether the CSS file is correctly linked to the expected HTML file.
@@
LLaMA-2 takes a different route compared to ChatGPT and the original answer, by exploring solutions involving exclusively the code given as input ($\theta$ = 0.72). The solution proposed is related to possible properties being overridden by others, which ChatGPT and Stack Overflow answers also briefly discussed. To address the issue, LLaMA-2 suggests replacing the usage of the property max-width for min-width, or the previous two properties for just break-point. These examples show the ability of ChatGPT and LLaMA-2 to address different subjects, with a high similarity when compared to human answers, leading us to observe a good level of reliability.
@@
Evaluating reliability relying exclusively on textual similarity might introduce biases in our results, as we briefly discuss when presenting the question addressed by LLaMA-2, which was not properly correct. The opposite scenario can also occur; for example, a user asks for instructions about exporting a fillable textbox PDF file using PowerBI. The accepted answer reports that the required action is not possible to be done [^17] https://stackoverflow.com/questions/74639660/. ChatGPT correctly informs that the required task is currently impossible to do. However, the LLM provides further alternatives that might support the user, resulting in a low textual similarity ($\theta$ = 0.10). We may conclude that textual similarity represents a good metric for reporting high similarity between the answers but not for reporting low similarity, as false negatives might take place. To address this threat, we also relied on LLMs’ judgments to compute the semantic similarity between answers and analyzed the level of agreement in these judgments. Overall, LLaMA-2 reported a more positive evaluation with only 7.4% of the answers with low similarity, whereas ChatGPT reported a frequency of 22.1%. This result goes the opposite of our manual classification, where the frequency of low similarity was around 44% of our sample.
@@
When computing the Krippendorff alpha between ChatGPT and LLaMA-2, we observe a coefficient of 0.02. Although we have used the LLMs with their default configurations, we observe that their ratings are not in good agreement, leading us to conclude that their ratings may not be reliable for the given data. When computing the Krippendorff alpha for multiple runs using ChatGPT, we observed more consistent results but still had a low coefficient (0.34). This result shows that the LLMs, even on different runs of the same model, do not present consistent and stable results regarding judging the semantic similarity of the answers. We believe that the temperature and advanced approaches might play a role in our findings. However, even in such conditions, we might expect some inconsistency when judging semantic similarity. Finally, we also observe the same low similarity when evaluating the agreement of the LLMs with our manual analysis (Krippendorff alpha 0.26 and -0.21, LLaMA and ChatGPT, respectively). As a result, we can conclude that possible conclusions we draw here based on this semantic similarity would be biased by the random behavior of the LLMs (see Section 6).
Sentiment Analysis
Regarding the sentiment analysis performed, we observe that most of the answers provided by Stack Overflow users and LLMs present a neutral sentiment associated (96%, 97%, and 79%, Stack Overflow, ChatGPT, and LLaMA, respectively). Unlike other web forums, Stack Overflow users are used to keeping a formal, neutral, and direct way of speaking, which could be misinterpreted as demanding (negative sentiment), especially when listing a sequence of steps to be followed. When comparing the answers, we observe that ChatGPT reports a more neutral sentiment compared to the original answers from Stack Overflow users (Wilcoxon test, p-value <0.01). Furthermore, ChatGPT did not report any answer with a negative sentiment, while 13 user answers were classified with negative sentiment. Such a result reinforces the ability of ChatGPT to adapt itself when answering questions, specifically in a tech context evaluated here.
Comparing LLaMA answers with those generated by ChatGPT and Stack Overflow users, we observe a significant statistical difference, revealing that LLaMA-2 report a less neutral sentiment (Wilcoxon test, p-value <0.01 and 0.01, respectively). However, unlike the previous answers, LLaMA-2 reports more answers with positive sentiment (20%). These results show that, in its current version, LLaMA-2 consistently adopts a positive-neutral sentiment when prompted by users.
@@
RQ1 answer: Checking the reliability based on textual similarity is statistically consistent. Both before and after ChatGPT’s release, the generated answers show a high degree of similarity to the accepted answers on Stack Overflow, as evidenced by the textual similarity (cosine metric). Although textual similarity seems to be a good metric for reporting answers with high similarity, this metric is not applicable for reporting low similarity, as false negatives might occur. Finally, the adopted semantic similarity metric based on LLMs classification also does not present an applicable metric for checking similarity among the answers. Although ChatGPT presents a more neutral sentiment when generating their answers, LLaMA-2 presents a more positive sentiment.
@@
4.2 RQ2: What specific domains pose challenges for ChatGPT and LLaMA-2 when generating reliable and accurate answers?
Figure 5: Distribution of question domains for high- and low-similarity answers generated by LLMs (ChatGPT and LLaMA-2).
@@
This research question focuses on understanding the associated challenges faced when generating these answers and the factors that might lead the LLMs to generate more accurate answers. Next, we discuss each aspect individually.
Challenges for LLMs
To report the challenges faced by LLMs, we manually analyzed the questions associated with high and low textual similarity (see Section 3.3.2). Regarding the domain of the questions classified with low textual similarity for LLMs (Figure 5), we reach the same conclusions for both LLMs. Frameworks and Libraries represent the most challenging domain for the LLMs evaluated here (39.6% and 34.9%, ChatGPT and LLaMA, respectively), followed by Programming Languages (20.8% and 23.4%). When looking at the different groups of questions with the low textual similarity associated with each LLM, we observe that they share 64% of the same questions, revealing that despite the structural differences between the models, they partially share the same limitations when addressing these domains. Furthermore, less than 20% of all generated answers by both models presented a similarity of less than 0.5 (14% and 16%, for ChatGPT and LLaMA, respectively). This small frequency shows that even when the LLMs fail to create a more similar answer, they do not deviate from the original direction of the accepted answers.
On the other hand, for the questions with high textual similarity, the most promising domains are Programming Languages (40.5% and 40.3%), followed by Frameworks and Libraries (23.7% and 23%). Although the Frameworks and Libraries domain is previously reported as a challenging domain for LLMs, we must discuss how the type of questions associated with each group plays an important role here. When users ask for specific, popular, and standard features of these dependencies, the LLMs achieve a high similarity; however, when the questions address general conventions, standard practices or common knowledge adopted by the community which may not be explicitly documented, or unpopular features about using these external dependencies, a low similarity is observed. For example, a user asks for support when failing to authenticate to an application using valid credentials. The solution accepted guides the user to hash the password when creating a new user in the application: [^18] https://stackoverflow.com/questions/75727923/
{quoting}
You need to hash the password before creating the user […] You will need to create the user again since you will create a user with a hashed password to let the authentication succeed.
In its generated answer, ChatGPT even warns about the user’s credential and object creation steps without mentioning the actual cause.
Further exploring the questions, we could observe some structural aspects that might bring some light to our observations. For example, when we look for external references linked to questions regarding framework and dependencies, we observe that the LLMs performed poorly for questions with more external references. While ChatGPT and LLaMA had to deal with 10 and 13 questions with external references when reporting low similarity, they had to deal with only 2 and 4 questions, respectively, when reporting higher similarity. For example, consider a user facing an issue when logging out in a VueJS app using Supertokens for authentication. Even after clicking the Logout button, the session data remained (userInfo), and the user still had an active session. Attached to the question, the user also reports a link holding a code snippet provided by SuperTokens.191919 Currently, the provided link is unavailable; however, for each code snippet reported in the documentation, SuperTokens reports an associated explanation. When answering the question, the respondent clarifies that the browser was not saving session cookies because the frontend was running on localhost while the API was on 127.0.0.1, and cross-domain requests require HTTPS for browsers to store cookies properly. Without these extra external details, the LLMs take a different route, mainly guiding the user to check the setup of the environment and sign-out options.
Although the advanced version of ChatGPT can access external links using the web tool (GPT-4-turbo), allowing the model to perform searches and open URLs when needed, the evaluated models in this study do not have such a capability. By linking external content, the questions provide a less self- contained context, and consequently, users are expected to read and follow these links for complete understanding. As a result, these questions are more challenging for LLMs to address, as the models must rely solely on the information explicitly provided in the prompt without access to any external supplementary content.
Another structural aspect observed is related to the level of detail provided in the questions. Although both groups provide code as supportive material, we observe that LLMs report higher similarity for questions supported not only by code but also by more supplementary data, like output, expected outcome, and error messages. In this context, for questions with low similarity, we observe ChatGPT and LLaMA dealing with only code for 79% and 74% of the questions, respectively, while for questions with high similarity, the rate was around 60% and 62%, respectively (information collected as previously reported in Table 2). These additional data provide more details about the problem, allowing the LLMs to have a better understanding of the target context.
We observe the same behavior regarding questions addressing programming language tasks when being supported exclusively by code. However, the difference was smaller than the previously discussed domain (for low similarity 78% and 85%, and for high similarity 72% and 76%, for ChatGPT and LLaMA). We believe that programming language questions are more direct as they involve only the aspects of the language itself. For frameworks and libraries, further aspects might be considered, like environment setup and versioning. However, further studies are required to better understand how these aspects could be better handled when prompting LLMs for related questions.
@@
From all performed analyses, the results show only a statistically significant correlation between the number of tags adopted for a question and the similarity for the answers generated post ChatGPT release (p-value <0.05). The correlation coefficient shows a negative relationship with $\rho=-0.11$, meaning that while one variable increases, the other might decrease. This result indicates that when ChatGPT is prompted to address problems related to one specific tag, it reports more reliable answers (high similarity). While one might expect that providing more information (i.e., multiple tags) would improve reliability, we believe that it can have the opposite effect. For example, the two questions discussed in Section 4.1 only adopt one single tag (java-8 and media-queries, respectively). We believe that by restricting the question to one tag, the LLM can focus its reasoning on a well-defined topic, reducing the risk of generating off-topic responses. However, this negative relationship reveals that ChatGPT might struggle when different contexts are present in the same question, as it has to arrange different topics to generate an answer. When a question includes multiple tags, especially with unrelated concepts, the LLM may struggle to establish a clear reasoning path, potentially affecting the reliability of its answers. For instance, when a user asks for support regarding an unresolved reference, they use three tags, which are all related to each other (android, android-livedata, and android-lifecycle) [^20] https://stackoverflow.com/questions/75449324/. For this question, ChatGPT generated an answer opposite from the one provided by the Stack Overflow user, achieving a low cosine similarity ($\theta$ = 0.37).